OpenStudy (mathmath333):

whats the trick to find remainder of 2^14/7

ganeshie8 (ganeshie8):

hint : `2^3/7` leaves the remainder `1`

ganeshie8 (ganeshie8):

\[\large 2^{14} = 2^{3*4 + 2} = 4*8^4 \equiv 4*1^4 \equiv 4 \]

ganeshie8 (ganeshie8):

thats one way ^^ heard of Fermat's little theorem ? :)

OpenStudy (mathmath333):

no

ganeshie8 (ganeshie8):

Okay, then just use the earlier method... did u get how/why 8 became 1 ?

ganeshie8 (ganeshie8):

\[\large 2^{14} = 2^{3*4 + 2} = 4*\color{red}{8}^4 \equiv 4*\color{red}{1}^4 \equiv 4\]

OpenStudy (mathmath333):

yes (7+1)/7 gives remainder 1

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

another way to look at :\[\large 8^4 = (7+1)^4 \] familiar with binomial expansion ?

OpenStudy (anonymous):

You might find this useful later in life http://vikhyat.net/blog/archives/2011/03/30/remainder-when-dividing-large-numbers/